On January 1, 2019 company A issues a $1,500,000 of 10% bonds due in 10 years with interest payable at year end annually. The current market rate of interest bonds of similar risk is 8%

A) What journal entry will company A record when issuing the bond

B) What will the buyers pay for this bond issue?
Answer #1

n = 10, i = 8%
a) PV of Principal = (1,500,000 x 0.4632) = 694,800 (Present Value Factor of 8% for 10th year = .04632, as per Present Value table)
b) PV of interest payments
(1,500,000 x 0.10) = 150,000 payment
(150,000 x 6.71008) = 1,006,512 (Present Value Factor of 8% for 10 years = .6.71008, as per annuity table)


Present value price of bonds = 694,800 + 1,006,512 = 1,701,312

A)   journal entry company A record when issuing the bond:

Debit Cash 1,701,312

Credit Bond Payable 1,500,000

Credit Premium on Bonds Payable 201,312

B) Buyers pay for this bond issue? = Present value price of bonds = 694,800 + 1,006,512 = $1,701,312 (as computed above)

If you like the answer, kindly give a ?.